Opened 7 years ago

Closed 3 years ago

Last modified 2 years ago

#446 closed flight processing (fixed)

EM10/02, flight day 138/2012, Milos

Reported by: knpa Owned by: knpa
Priority: alpha 4 medium Milestone: 2012 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by adbe)

Data location: /users/rsg/arsf/arsf_data/2012/flight_data/greece/EM10_02-2012_138_Milos

Data arrived from Greece via SATA disk on 13/07/2012

EM10/02 Scientific objective: Study of a volcano on the island of Milos, Greece.

EM10/02 Priority: a4m

PI: Graham Ferrier

Sensors:

Camera (17/08/2012)
Eagle (30/08/2012)
Hawk (30/08/2012)
LiDAR (17/08/2012)

Change History (17)

comment:1 Changed 7 years ago by knpa

Eagle raw names were a confusing jumble (e.g. VNIR138-12-1-1-2.raw). Have renamed as 1 to 15 to match the Hawk.

There was a crash on eagle line 11 and it is split into two parts. The first part has no log file.

comment:2 Changed 7 years ago by adbe

Processed Navigation.

comment:3 Changed 7 years ago by adbe

Deliveries created for Camera and LiDAR

comment:4 Changed 7 years ago by chho

Camera delivery check no problems, ready to deliver

comment:5 Changed 7 years ago by anhi

Lidar delivery check

  • incorrect Northing value in the Readme
  • Spelling mistake in the Readme

All corrected and ready to deliver

comment:6 Changed 7 years ago by adbe

LiDAR:

  • Corrected Northing value.
  • Corrected spelling mistake

comment:7 Changed 7 years ago by adbe

  • Description modified (diff)

Camera and LiDAR delivered to Graham Ferrier on 17/08/2012 on disk 125.

comment:8 Changed 6 years ago by adbe

Hyperspectral delivery created.

comment:9 Changed 6 years ago by anhi

Hyperspectral delivery check complete

  • The logsheet eagle file numbers are wrong, correct in the Readme
  • Dropped scans in eagle and hawk lines 5 and 13
  • Zipped map files need creating

comment:10 Changed 6 years ago by adbe

Made changes to hyperspectral

comment:11 Changed 6 years ago by adbe

  • Description modified (diff)

Delivered Hyperspectral to Graham Ferrier on disk 127 on 30/08/2012

comment:12 Changed 6 years ago by knpa

  • Component changed from Processing: general to Archiving
  • Owner changed from adbe to knpa
  • Status changed from new to assigned

Beginning archiving.

comment:13 Changed 4 years ago by asm

Archiving
started archiving

comment:14 Changed 3 years ago by dac

Archiving

Uploading data to NEODC

comment:15 Changed 3 years ago by dac

  • Resolution set to fixed
  • Status changed from assigned to closed

Archiving

Raw and processed data available from NEODC:

http://browse.ceda.ac.uk/browse/neodc/arsf/2012/EM10_02/EM10_02-2012_138_Milos

Marking as closed.

comment:16 Changed 2 years ago by asm

Data Delivery

Upon request, data has been dispatched via hard drive on 17th January 2017 to address in Nottingham of PhD. (from School of Earth Sciences; University of Bristol).

comment:17 Changed 2 years ago by asm

Data Delivery

The above described user confirmed reception on 20th January 2017.

Last edited 2 years ago by asm (previous) (diff)
Note: See TracTickets for help on using tickets.